# Deterministic Dynamic Edge-Colouring

### v1

@@ 3.2 Algorithms for constructing and extending Vizing chains
Similarly to before, we can perform the above extension in $\tilde{O}(\Delta_{\max}^{3}+\Delta_{\max}\cdot{}t)$ time. We will say that we extend the Vizing chain via Lemma 13 in a consistent manner. Note that if every extension in a Vizing chain $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ is consistent, then $P_{j+1}$ will never overlap $P_{j}$ in any edges for all $j$. Indeed, the colours of $P_{j+1}$ are either disjoint from those of $P_{j}$ or identical to those of $P_{j}$. In the latter case, $P_{j+1}$ and $P_{j}$ belong to different components in the shifted colouring.
@@
Definition 14.
Let $VC=F_{1}+P_{1}+\dots+F_{j}+P_{j}$ be a non-overlapping multi-step Vizing chain. We say that $VC$ is consistent if the following conditions hold:
  • $F_{1}+P_{1}$ is constructed in a consistent manner (Lemma 12 and following discussion).
  • For $i\geq 2$ we have that $F_{i}+P_{i}$ is constructed in a consistent manner (Lemma 13 and following discussion).
[…]
@@ 3.3 Path-stepping processes
We will now consider a looser process than the one used to construct Vizing chains. Beginning from some vertex $w$, we let one step be the process of first moving to a neighbour of $w$, say $y$, and then following some bichromatic path $P$ having $y$ as an endpoint. As with the multi-step Vizing chains we are allowed to truncate the bichromatic path at any point to end up at some new vertex $w^{\prime}$. We may then iteratively repeat this process at $w^{\prime}$. If we do so $i$ times, we say that we run an $i$-step process starting from $w_{1}$. We denote by $w_{1}+y_{1}+\hat{P}_{1}+w_{2}+y_{2}+\hat{P}_{2}+\dots w_{i}+y_{i}+\hat{P}_{i}$ the $i$-step process beginning at $w$, that first jumps to the neighbour $y_{1}$ of $w$, then follows a truncated maximal bichromatic path $P_{1}$ to end up at $w_{2}$, and then jumps to the neighbour $y_{2}$ of $w_{2}$ and follows a truncated maximal bichromatic path $P_{2}$ and so on and so forth. We will say that the process is non-overlapping if for all $k,j$ we have that $V(P_{k})\cap V(P_{j})\neq\emptyset$ implies that $|k-j|\leq 1$, if $E(P_{k})\cap E(P_{j})\neq\emptyset$ implies $k=j$, if $V(P_{k})\cap N^{2}(w_{j})\neq\emptyset$ for $k\geq j$ implies that $k=j$, and, finally, if the maximal bichromatic path $Q_{j}$ containing $P_{j}$ is equal to the maximal bichromatic path $Q_{k}$ containing $P_{k}$, then $j=k$. In other words, we will only allow a bichromatic path to intersect the 2-hop-neighbourhood of the immediately preceding path, and we will not allow it to intersect the edge set of any previous paths. If $|P_{j}|+1\leq\ell$ for all $j$, we say that the process has step length at most $\ell$. If $P_{i}$ is a maximal bichromatic path, we say that the $i$-step process is augmenting.
@@
Maintaining information about this type of process is useful, as it allows us to accurately represent where consistent Vizing chains might go, while at the same time avoiding the need to worry about exactly which extensions are possible at what times. In particular, we observe that any consistent and non-overlapping $i$-step Vizing chain $F_{1}+P_{1}+\dots+F_{i}+P_{i}$ where $F_{1}$ is centered at $w$ naturally induces an $i$-step process on $w$, by letting $w_{j}=\text{center}(F_{j})$, letting $y_{j}$ be the second point of $P_{j}$, and letting $\hat{P}_{j}$ be the the bichromatic sub-path of the path $P_{j}$ excluding only the first edge of the chain $P_{j}$. To make this observation more precise, we will define and consider semi-consistent processes. Here, we will put some further restrictions on the processes so that they model consistent Vizing chains exactly.
@@
Given an $i$-step process $w_{1}+y_{1}+P_{1}+w_{2}+y_{2}+P_{2}+\dots w_{i}+y_{i}+P_{i}$, we say that the extension $w_{j}+y_{j}+P_{j}$ for $j\geq 2$ is consistent with Lemma 13, if the following is the case: consider the colouring $c^{\prime}$ obtained from $c$ as follows: suppose $P_{j-1}=p_{1}p_{2}\dots p_{t}$ is $(\kappa_{1},\kappa_{2})$-bichromatic with $c(p_{t-1}p_{t})=\kappa_{1}$. Suppose the $P=p_{1}p_{2}\dots p_{t}p_{t+1}\dots p_{s}$ is a maximal bichromatic path containing $P_{j-1}$. Note that if $s=t$, then $P_{j-1}$ is augmenting, and it cannot be extended consistently.
@@
We obtain $c^{\prime}$ by first un-colouring the edge $p_{t}p_{t+1}$ i.e. $c^{\prime}(p_{t}p_{t+1})=\neg$, and then setting $c^{\prime}(p_{k}p_{k+1})=\kappa_{2}$ if $c(p_{k}p_{k+1})=\kappa_{1}$ and vice versa for $k<t-1$ setting $c^{\prime}(p_{k}p_{k+1})=\kappa_{1}$ if $c(p_{k}p_{k+1})=\kappa_{2}$. It is straight-forward to check that if $c$ is a proper $(\Delta_{\max}+1)$-edge-colouring, then so is $c^{\prime}$. Now, apply Lemma 13 to $p_{t-1}p_{t}$ with the colouring $c^{\prime}$, $u=p_{t-1}$, $v=p_{t}$, and the colours $\kappa_{1}$ and $\kappa_{2}$ as before. By fixing a length $t^{\prime}=|P_{i}|$, the lemma together with the discussion that follows it produces a fan chain $F$ and a bichromatic-path chain $Q=(q_{0}q_{1},\dots,q_{\hat{t}}q_{\hat{t}+1})$ with $\hat{t}\leq t^{\prime}$ such that $F+Q$ is a $1$-step Vizing chain. We say the extension $w_{j}+y_{j}+P_{j}$ for $j\geq 2$ is consistent with Lemma 13 if $w_{j}=\text{center}(F_{j})=q_{0}$, and if $y_{j}=q_{1}$, and finally $P_{j}=q_{1},q_{2},\dots,q_{t^{\prime}}$. If $\hat{t}=t^{\prime}-1$, then $P_{j}$ is augmenting. If $\hat{t}<t^{\prime}-1$, then the extension is not consistent.
@@
  • For $j\geq 2$ we have that $w_{j},y_{j}$ and $P_{j}$ is an extension consistent with Lemma 13.
@@
We observe that for any $j<i$ the length of $P_{j}$ is at least $3$. Indeed, otherwise Item 2 in Definition 15 fails. The following Lemma is immediate:
Lemma 16.
Let $G$ be a graph endowed with a proper (partial) edge-colouring $c$. Let $\mathcal{SCP}$ be the set of semi-consistent processes on $(G,c)$, and let $\mathcal{NOP}$ be the set of non-overlapping processes on $(G,c)$. Then we have:
$\mathcal{SCP}\subset\mathcal{NOP}$
Semi-consistent processes are in correspondence with consistent Vizing chains in the following sense:
Lemma 17.
There is a mapping $\gamma$ from the set of consistent $i$-step Vizing chains to the set of semi-consistent $i$-step processes. In particular, if a consistent $i$-step Vizing chains augmenting, so is its image, and similarly if a semi-consistent $i$-step process is augmenting, so is every Vizing chain in its pre-image.
Proof.
Let $VC=F_{1}+Q_{1}+\dots+F_{i}+Q_{i}$ be any consistent $i$-step Vizing chain. Then we will let $VC=F_{1}+Q_{1}+\dots+F_{i}+Q_{i}$ correspond to the $i$-step process $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$, where $w_{j}=\text{center}(F_{j})$, if $Q_{j}=(q_{0}q_{1},\dots,q_{t_{j}+1})$, then $y_{j}=q_{1}$ and $P_{j}=q_{1},q_{2}\dots,q_{t_{j}}$.
Next we show that $C$ is semi-consistent. First of all, we need to check that $C$ is non-overlapping. This follows from the fact that $VC$ is consistent, which puts precisely the same constraints on the $P_{j}$’s due Definition 14 as is required for $C$ to be non-overlapping. The first item of Definition 15 is satisfied since the second item of Definition 14 holds. This is the case since no edge outside of $P_{j-1}$ in the two-hop neighbourhood of the center of $F_{j}$ has changed colour before Lemma 13 is applied. This means that applying it directly to $P_{j-1}$ yields the same result as using it to extend the Vizing chain. Finally, items 3 and 4 of Definition 14 are equivalent to the last two items of Definition 15.
@@
The other direction is symmetrical. ∎
@@
Lemma 18.
Let $G$ be a graph with maximum degree $\Delta\leq\Delta_{\max}$, and let $c$ be a proper $(\Delta_{\max}+1)$-edge-colouring of $G$. Let $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ be an augmenting semi-consistent $i$-step process with step-length $\leq\ell$. Let $e=w_{1}v$ be an un-coloured edge. Suppose that applying Lemma 12 and the following discussion yields a consistent $1$-step Vizing chain $(F+Q_{1})_{\ell}$ such that
[…]
  • $y_{1}=x_{k}$ with $F=(w_{1}x_{1},w_{1}x_{2},\dots w_{1}x_{k}$
  • $P_{1}=q_{1}q_{2}\dots q_{\ell}$ with $Q=(q_{0}q_{1},q_{1}q_{2},\dots,q_{\ell}q_{\ell+1})$
@@
Proof.
We will construct the augmenting consistent Vizing chain on $e$, say $VC=F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{i}+Q_{i}$ as follows: we set $F_{1}=F$ and $Q_{1}=Q$. Then we extend $F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j}$ by applying Lemma 13 and the following discussion making sure that the length of $F_{j+1}+Q_{j}$ is equal to $t_{j}+2$ with $t_{j}=|P_{j}|$.
We now observe that $\gamma(VC)=C$. Indeed, by construction every extension results in $Q_{j}=(q_{0,j}q_{1,j},\dots,q_{t_{j}+1,j}q_{t_{j}+2,j})$ such that $P_{j}=q_{1,j}q_{2,j}\dots q_{t_{j}+1,j}$. This follows from the fact that since $C$ is semi-consistent, the extension performed on $P_{j}$ uses a colouring $c^{\prime}$ which only differs from the colouring $\operatorname{Shift}(c,F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j})$ outside of the $2$-hop neighbourhood of $w_{j}$. Indeed, $c^{\prime}$ and $\operatorname{Shift}(c,F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j})$ agree on the colours along $P_{j}$ by construction, and $w_{j}+1$ is disjoint from the two-hop neighbourhood of $C=w_{1}+y_{1}+P_{1}+\dots+w_{j-1}+y_{j-1}+P_{j-1}+w_{j}$ also by construction. Since $C$ is semi-consistent, the extension ensures that $Q_{j}=(q_{0,j}q_{1,j},\dots,q_{t_{j}+1,j}q_{t_{j}+2,j})$ such that $P_{j}=q_{1,j}q_{2,j}\dots q_{t_{j}+1,j}$.
It now follows that $\gamma(VC)=C$, and therefore Lemma 17 now guarantees that $VC$ is also augmenting. ∎
@@ 4 Dynamic t-splitters
Given a graph $G$ equipped with any (not necessarily proper) edge-colouring $c$, we let $d_{i}(u)$ be the degree of $u$ in the graph induced in $G$ by all edges coloured with the colour $i$. We let $\tilde{d}=\max_{\kappa}\max_{v}d_{\kappa}(v)$. Given two disjoint subsets of vertices $A,B\subset V(G)$, we let $G[A,B]$, be the bipartite subgraph of $G$ on vertex set $A\cup B$ containing exactly the edges going from a vertex in $A$ to a vertex in $B$. We let $s_{i}(v)=\max\{d_{i}(v)-\frac{\Delta_{\max}}{t},0\}$ be the surplus of $i$-coloured edges around $v$. Finally, we let $L:E(G)\rightarrow\mathbb{Z}$ be a function that maps $uv\mapsto s_{c(uv)}(u)+s_{c(uv)}(v)$, and we set $\max L(e)=\tilde{L}$.
In this section, we will show Theorem 3, which we restate below for convenience:
Theorem 19 (Identical to Theorem 3).
Let $G$ be a dynamic graph, and suppose $\varepsilon>0$, $t\in\mathbb{N}$, and an upper bound $\Delta_{\max}$ on the maximum degree of $G$ through the entire update sequence is given. Assume that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then there is a dynamic algorithm that computes a $t$-splitter of $G$ with an amortised update-time of $O(t^{3}\varepsilon^{-4}\log^{3}m)$.
@@
We first show that if $1\leq\eta\leq\tfrac{\varepsilon^{2}}{128\log n}\tfrac{\Delta_{\max}}{t}$, then $\tilde{d}\leq(1+\varepsilon)\frac{\Delta_{\max}}{t}$. We note already now that we will only apply the above Theorem, when $\tfrac{\Delta_{\max}}{t}=\Omega(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$, so the assumption that $\eta\geq 1$ is not dangerous.

### v2

@@ 3.2 Algorithms for constructing and extending Vizing chains
Similarly to before, we can perform the above extension in $\tilde{O}(\Delta_{\max}^{3}+\Delta_{\max}\cdot{}t)$ time. We will say that we extend the Vizing chain via Lemma 3.5 in a consistent manner. Note that if every extension in a Vizing chain $F_{1}+P_{1}+F_{2}+P_{2}+\dots F_{i}+P_{i}$ is consistent, then $P_{j+1}$ will never overlap $P_{j}$ in any edges for all $j$ (except possibly in the first edge). Indeed, the colours of $P_{j+1}$ are either disjoint from those of $P_{j}$ or identical to those of $P_{j}$. In the latter case, $P_{j+1}$ and $P_{j}$ belong to different components in the shifted colouring.
@@
Definition 3.6.
Let $VC=F_{1}+P_{1}+\dots+F_{j}+P_{j}$ be a multi-step Vizing chain. We say that $VC$ is consistent if the following conditions hold:
  • $F_{1}+P_{1}$ is constructed in a consistent manner (Lemma 3.4 and following discussion).
  • For $i\geq 2$ we have that $F_{i}+P_{i}$ is constructed in a consistent manner (Lemma 3.5 and following discussion).
  • $\text{center}(F_{j})$ is disjoint from the $2$-hop neighbourhood of $\text{center}(F_{1})+P_{1}+\dots+\text{center}(F_{j-2})+P_{j-2}+\text{center}(F_{j-1})$ for all $j$.
  • For all $j$, we have that $P_{j}$ is disjoint from the $2$-hop neighbourhood of $\text{center}(F_{k})$ for all $k\leq j-1$ and from $P_{k}$ for all $k\leq j-2$.
@@ 3.3 Path-stepping processes
We will now consider a looser process than the one used to construct Vizing chains. Beginning from some vertex $w$, we let one step be the process of first moving to a neighbour of $w$, say $y$, and then following some bichromatic path $P$ having $y$ as an endpoint. As with the multi-step Vizing chains, we are allowed to truncate the bichromatic path at any point to end up at some new vertex $w^{\prime}$. We may then iteratively repeat this process at $w^{\prime}$. If we do so $i$ times, we say that we run an $i$-step process starting from $w_{1}$. We denote by $w_{1}+y_{1}+\hat{P}_{1}+w_{2}+y_{2}+\hat{P}_{2}+\dots w_{i}+y_{i}+\hat{P}_{i}$ the $i$-step process beginning at $w_{1}$, that first jumps to the neighbour $y_{1}$ of $w_{1}$, then follows a truncated maximal bichromatic path $P_{1}$ to end up at $w_{2}$, and then jumps to the neighbour $y_{2}$ of $w_{2}$ and follows a truncated maximal bichromatic path $P_{2}$ and so on and so forth. We will say that the process is non-overlapping if for all $k,j$ we have that $V(P_{k})\cap V(P_{j})\neq\emptyset$ implies that $|k-j|\leq 1$, if $E(P_{k})\cap E(P_{j})\neq\emptyset$ implies $k=j$, if $V(P_{k})\cap N^{2}(w_{j})\neq\emptyset$ for $k\geq j$ implies that $k=j$, and, finally, if the maximal bichromatic path $Q_{j}$ containing $P_{j}$ is equal to the maximal bichromatic path $Q_{k}$ containing $P_{k}$, then $j=k$. In other words, we will only allow a bichromatic path to intersect the 2-hop-neighbourhood of the immediately preceding path, and we will not allow it to intersect the edge set of any previous paths. If $|V(P_{j})|\leq\ell$ for all $j$, we say that the process has step length at most $\ell$. If $P_{i}$ is a maximal bichromatic path, we say that the $i$-step process is augmenting.
@@
Maintaining information about this type of process is useful, as it allows us to accurately represent where consistent Vizing chains might go, while at the same time avoiding the need to worry about exactly which extensions are possible at what times. In particular, we observe that any consistent and non-overlapping $i$-step Vizing chain $F_{1}+P_{1}+\dots+F_{i}+P_{i}$ where $F_{1}$ is centered at $w$ naturally induces an $i$-step process on $w$, by letting $w_{j}=\text{center}(F_{j})$, letting $y_{j}$ be the second point of $P_{j}$, and letting $\hat{P}_{j}$ be the the bichromatic sub-path of the path $P_{j}$ excluding only the first and the last edge of the chain $P_{j}$. To make this observation more precise, we will define and consider semi-consistent processes. Here, we will put some further restrictions on the processes so that they model consistent Vizing chains exactly.
@@
Given an $i$-step process $w_{1}+y_{1}+P_{1}+w_{2}+y_{2}+P_{2}+\dots w_{i}+y_{i}+P_{i}$, we say that the extension $w_{j}+y_{j}+P_{j}$ for $j\geq 2$ is consistent with Lemma 3.5, if the following is the case: consider the colouring $c^{\prime}$ obtained from $c$ as follows: suppose $P_{j-1}=p_{1}p_{2}\dots p_{t}$ is $(\kappa_{1},\kappa_{2})$-bichromatic with $c(p_{t-1}p_{t})=\kappa_{1}$. Suppose $P=p_{1}p_{2}\dots p_{t}p_{t+1}\dots p_{s}$ is a maximal bichromatic path containing $P_{j-1}$. Note that if $s=t$, then $P_{j-1}$ is augmenting, and it cannot be extended consistently.
@@
We obtain $c^{\prime}$ by first un-colouring the edge $p_{t}p_{t+1}$ i.e. $c^{\prime}(p_{t}p_{t+1})=\neg$, and then for all k<t1k<t-1 setting $c^{\prime}(p_{k}p_{k+1})=\kappa_{2}$ if $c(p_{k}p_{k+1})=\kappa_{1}$ and vice versa setting $c^{\prime}(p_{k}p_{k+1})=\kappa_{1}$ if $c(p_{k}p_{k+1})=\kappa_{2}$. It is straight-forward to check that if $c$ is a proper $(\Delta_{\max}+1)$-edge-colouring, then so is $c^{\prime}$. Now, apply Lemma 3.5 to $p_{t}p_{t+1}$ with the colouring $c^{\prime}$, $u=p_{t}$, $v=p_{t+1}$, and the colours $\kappa_{1}$ and $\kappa_{2}$ as before. By fixing a length $t^{\prime}=|P_{i}|$, the lemma together with the discussion that follows it produces a unique fan chain $F$ and a bichromatic-path chain $Q=(q_{0}q_{1},\dots,q_{\hat{t}}q_{\hat{t}+1})$ with $\hat{t}\leq t^{\prime}$ such that $F+Q$ is a $1$-step Vizing chain. We say the extension $w_{j}+y_{j}+P_{j}$ for $j\geq 2$ is consistent with Lemma 3.5 if $w_{j}=\text{center}(F_{j})=q_{0}$, and if $y_{j}=q_{1}$, and finally $P_{j}=q_{1},q_{2},\dots,q_{\hat{t}+1}$ if $\hat{t}+1\leq t^{\prime}$ and $P_{j}=q_{1},q_{2},\dots,q_{\hat{t}}$ otherwise. Note that in the former case $P_{j}$ is augmenting, and that in all cases the extension is unique.
Remark 3.7.
It is important to remark that in the case where $P_{j}$ is augmenting, we must have that $i=j$, and perhaps more importantly, that it could be the case that $F_{j}$ is not a maximal fan.
@@
  • For $j\geq 2$ we have that $w_{j},y_{j}$ and $P_{j}$ is an extension consistent with Lemma 3.5.
@@
We observe that for any $j<i$ the length of $P_{j}$ is at least $3$. Indeed, otherwise Item 2 in Definition 3.8 fails. Semi-consistent processes are in correspondence with consistent Vizing chains in the following sense:
Lemma 3.9.
There is a mapping $\gamma$ from the set of consistent $i$-step Vizing chains to the set of semi-consistent $i$-step processes. In particular, if a consistent $i$-step Vizing chain is augmenting, so is its image, and similarly if a semi-consistent $i$-step process is augmenting, so is every Vizing chain in its pre-image.
Proof 3.
Let $VC=F_{1}+Q_{1}+\dots+F_{i}+Q_{i}$ be any consistent $i$-step Vizing chain. Then we will let $VC=F_{1}+Q_{1}+\dots+F_{i}+Q_{i}$ map to the $i$-step process $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$, where for all $j$ we have $w_{j}=\text{center}(F_{j})$, $y_{j}=q_{1}$, and $P_{j}=q_{1},q_{2}\dots,q_{t_{j}}$ with $Q_{j}=(q_{0}q_{1},\dots,q_{t_{j}+1})$. If $VC$ is not augmenting, we let $w_{i}=\text{center}(F_{i})$, $y_{i}=q_{1}$, and $P_{i}=q_{1},q_{2}\dots,q_{t_{i}}$, otherwise we set $w_{i}=\text{center}(F_{i})$, $y_{i}=q_{1}$, and $P_{i}=q_{1},q_{2}\dots,q_{t_{i}+1}$.
Next we show that $C$ is semi-consistent. First of all, we need to check that $C$ is non-overlapping. This follows from the fact that $VC$ is consistent, which puts precisely the same constraints on the $P_{j}$’s in Definition 3.6 as is required for $C$ to be non-overlapping. The first item of Definition 3.8 is satisfied since the second item of Definition 3.6 holds. This is the case since no edge outside of $P_{j-1}$ in the two-hop neighbourhood of the center of $F_{j}$ has changed colour before Lemma 3.5 is applied. This means that applying it directly to $P_{j-1}$ yields the same fan and path choice as using it to extend the Vizing chain. Finally, items 3 and 4 of Definition 3.6 are equivalent to the last two items of Definition 3.8.
@@
The other direction is symmetrical.
@@
Lemma 3.10.
Let $G$ be a graph with maximum degree $\Delta\leq\Delta_{\max}$, and let $c$ be a proper $(\Delta_{\max}+1)$-edge-colouring of $G$. Let $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$ be an augmenting semi-consistent $i$-step process with step-length $\leq\ell$. Let $e=w_{1}v$ be an un-coloured edge. Suppose that applying Lemma 3.4 with $t=\text{length}(P_{1})+1$ and the following discussion yields a consistent $1$-step Vizing chain $(F+Q_{1})_{t}$ such that
  • $\text{center}(F)=w_{1}$.
  • $y_{1}=x_{k}$ with $F=(w_{1}x_{1},w_{1}x_{2},\dots w_{1}x_{k})$.
  • $Q=(q_{0}q_{1},q_{1}q_{2},\dots,q_{t-1}q_{t})$ where $P_{1}=q_{1}q_{2}\dots q_{t-1}$.
@@
Proof 4.
We will construct the augmenting consistent Vizing chain on $e$, say $VC=F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{i}+Q_{i}$ as follows: we set $F_{1}=F$ and $Q_{1}=Q$. Then we extend $F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j}$ by applying Lemma 3.5 and the following discussion making sure that the length of $F_{j+1}+Q_{j}$ is equal to $t_{j}+1$ with $t_{j}=|P_{j}|+1$.
We now observe that $\gamma(VC)=C$. Indeed, by construction every extension results in $Q_{j}=(q_{0,j}q_{1,j},\dots,q_{t_{j}+1,j}q_{t_{j},j})$ such that $P_{j}=q_{1,j}q_{2,j}\dots q_{t_{j}-1,j}$. This follows from the fact that since $C$ is semi-consistent, the extension performed on $P_{j}$ uses a colouring $c^{\prime}$ which only differs from the colouring $\operatorname{Shift}(c,F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j})$ outside of the $2$-hop neighbourhood of $w_{j}$. Indeed, $c^{\prime}$ and $\operatorname{Shift}(c,F_{1}+Q_{1}+F_{2}+Q_{2}+\dots+F_{j}+Q_{j})$ agree on the colours along $P_{j}$ by construction, and $w_{j+1}$ is disjoint from the two-hop neighbourhood of $C=w_{1}+y_{1}+P_{1}+\dots+w_{j-1}+y_{j-1}+P_{j-1}+w_{j}$ also by construction. Since $C$ is semi-consistent, the extension ensures that $Q_{j}=(q_{0,j}q_{1,j},\dots,q_{t_{j}+1,j}q_{t_{j},j})$ where $P_{j}=q_{1,j}q_{2,j}\dots q_{t_{j}-1,j}$.
It now follows that $\gamma(VC)=C$, and therefore Lemma 3.9 now guarantees that $VC$ is also augmenting.
@@ 4 Dynamic t-splitters
Given a graph $G$ equipped with any (not necessarily proper) edge-colouring $c$, we let $d_{i}(u)$ be the degree of $u$ in the graph induced in $G$ by all edges coloured with the colour $i$. We let $\tilde{d}=\max_{\kappa}\max_{v}d_{\kappa}(v)$. Given two disjoint subsets of vertices $A,B\subset V(G)$, we let $G[A,B]$, be the bipartite subgraph of $G$ on vertex set $A\cup B$ containing exactly the edges going from a vertex in $A$ to a vertex in $B$. We let $s_{i}(v)=\max\{d_{i}(v)-\frac{\Delta_{\max}}{t},0\}$ be the surplus of $i$-coloured edges around $v$. Finally, we let $L:E(G)\rightarrow\mathbb{Z}$ be a function that maps $uv\mapsto s_{c(uv)}(u)+s_{c(uv)}(v)$, and we set $\max\limits_{e\in E}L(e)=\tilde{L}$.
In this section, we will show Theorem 1.3, which we restate below for convenience:
Theorem 4.1 (Identical to Theorem 1.3).
Let $G$ be a dynamic graph, and suppose $\varepsilon>0$, $t\in\mathbb{N}$, and an upper bound $\Delta_{\max}$ on the maximum degree of $G$ throughout the entire update sequence is given. Assume that $\tfrac{\Delta_{\max}}{t}\geq\tfrac{10^{4}\log^{2}n}{\varepsilon^{2}}$. Then there is a dynamic algorithm that computes a $t$-splitter of $G$ with an amortised update-time in $O(t^{3}\varepsilon^{-4}\log^{3}m)$ and an amortised recourse in $O(\tfrac{\log n}{\varepsilon^{2}})$.
@@
We first show that if $1\leq\eta\leq\tfrac{\varepsilon^{2}}{128\log n}\tfrac{\Delta_{\max}}{t}$, then $\tilde{d}\leq(1+\varepsilon)\frac{\Delta_{\max}}{t}$. We note already now that we will only apply the above Theorem, when $\tfrac{\Delta_{\max}}{t}=\Omega(\operatorname{\text{{\rm poly}}}(\varepsilon^{-1},\log n))$, so the assumption that $\eta\geq 16$ can always be realised when necessary.